OpenStudy (anonymous):

The following function defines a recursive sequence. f(0)=-4 f(1)=12 f(n)=-3xf(n-1)-2xf(n-2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, ... -4, -12, -28, -60, ... -4, 12, -18, 54, ... -4, 12, -18, -54, ...

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

Please help! D:

OpenStudy (anonymous):

@cwrw238 Can you help me?

OpenStudy (anonymous):

The notation is a little odd but the problem actually isn't that bad.

OpenStudy (anonymous):

n represents the term of the function that you want to find.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So, for n = 2, that means it's saying \[f(2) = -3(f(1)) - 2(f(0))\]

OpenStudy (anonymous):

We know both of those values (they're given) so we can just plug in. \[f(2) = -3(12) - 2(-4)\]\[f(2) = -28\]

OpenStudy (anonymous):

Now we know f(2), which means we can evaluate f(3), because f(3) is just \[f(3) = -3(f(2)) - 2(f(1))\]

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok. :D

OpenStudy (anonymous):

:D Can you walk me through the evaluation real quick though? ._.

OpenStudy (anonymous):

\[f(3) = -3(f(2) - 2(f(1))\]\[f(3) = -3(-28) - 2(12)\] Simplify that. :)

OpenStudy (anonymous):

Ah, thank you so much!

OpenStudy (anonymous):

Okay, so I got 84 - 24 or, 60. What do I do with that?

OpenStudy (anonymous):

That's the next entry in the sequence.

OpenStudy (anonymous):

So far the sequence we've generated is -4, 12, -28, 60.

OpenStudy (anonymous):

Is that all I need for the answer?

OpenStudy (anonymous):

You tell me. :P

OpenStudy (anonymous):

I think so.